OpenStudy (anonymous):

the sum of three consecutive numbers is -27. what are the numbers {-11, -13, -15} {-7, -9, -11} {-10, -9, -8} {-5, -9, -13}

OpenStudy (anonymous):

@Skrteljnr @texaschic101

OpenStudy (texaschic101):

do you know what consecutive numbers are ?

OpenStudy (anonymous):

they are numbers that follow eachother

OpenStudy (anonymous):

I wanna say C

OpenStudy (texaschic101):

consecutive numbers are numbers one right after the other....for example : 1,2,3,4,5...these are consecutive. So lets let one number be x, the other number is x + 1. and the last number is x + 2. added together, they equal -27. x + x + 1 + x + 2 = -27 -- combine like terms 3x + 3 = -27 -- subtract 3 from both sides 3x = - 27 - 3 3x = -30 -- divide by 3 x = - 10 x + 1 = -10 + 1 = -9 x + 2 = -10 + 2 = -8 numbers are -10,-9,-8 and that would be C....you are correct :)

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :)